OpenStudy (anonymous):

Identify the class (or method to use) for the equation and solve #1) sin(x+y)dx + [2y+sin(x+y)]dy=0 #2) dy/dx=(sinx)y+2xe^(-cosx) Thank you in advance for the help! Really stuck on these two!

OpenStudy (anonymous):

1) might be an exact equation. Not sure.

OpenStudy (anonymous):

I know the method and the solutions, I just don't know how to solve it.. #1 is exact and the general solution is y^2-cos(x+y)=constant #2 is linear with the general solution y=[e^(-cosx)](x^2+C) If anyone could show me the steps to get these answers, I would greatly appreciate it!! thank you!

ganeshie8 (ganeshie8):

#1) sin(x+y)dx + [2y+sin(x+y)]dy=0 compare wid below : \(\large \mathbb{f_x dx + f_y dy = 0}\) you need to find some function \(\large \mathbb{ f(x,y)}\), such that : \(\large \mathbb{f_x = \sin(x+y)}\) \(\large \mathbb{f_y = 2y+\sin(x+y)}\)